OpenStudy (burak00b):

Question: A poorly-run restaurant messes up one order in three. You and two people from work each eat there once this month, each of you eating there on a different day. What is the probability that this month the restaurant didn’t mess up any of the orders placed by your or your two friends? (Your answer should be a fraction. Assume independence.)

OpenStudy (mathmale):

Think of the problem this way: P(messed up order) = 1/3 P(order not messed up) = 1 - 1/3 = 2/3 that's for one person on one day. Using this info, how would you address this question? What is the joint probability of independent events?

OpenStudy (mathmale):

suppose A and B are independent events, and P(A) and P(B) are known. What is the probability that these events will occur simultaneously? "Joint probability."